# Speed, Accuracy, and Complexity

### v1

@@ Abstract
This paper re-examines the validity of using response time to infer problem complexity. It revisits a canonical Wald model of optimal stopping, taking signal-to-noise ratio as a measure of problem complexity. While choice quality is monotone in problem complexity, expected stopping time is inverse $U$-shaped. Indeed decisions are fast in both very simple and very complex problems: in simple problems it is quick to understand which alternative is best, while in complex problems it would be too costly — an insight which extends to general costly information acquisition models. This non-monotonicity also underlies an ambiguous relationship between response time and ability, whereby higher ability entails slower decisions in very complex problems, but faster decisions in simple problems. Finally, this paper proposes a new method to correctly infer problem complexity based on the finding that choices react more to changes in incentives in more complex problems. Keywords: Sequential Sampling; Complexity; Information Acquisition; Drift-Diffusion; Statistical Decision Theory. JEL Classifications: D83, D84, D90.
@@ 1.  Introduction
It is almost definitional that more complex decision problems entail more mistakes. Individuals make more mistakes and choose dominated alternatives in more complex environments, both in experimental settings (e.g. RabinWeizsacker2009AER; Martinez-MarquinaNiederleVespa2019AER; BanovetzOprea2023AEJMicro; Enke2020QJE) and in real-world decisions with significant financial implications — e.g. selecting health insurance plans (SinaikoHirth2011JHE; BhargavaLoewensteinSydnor2017QJE), pension plans (ChoiLaibsonMadrian2011REStat), and mortgages (KeysPopePope2016JFE; AgarwalRosenYao2016MnSc). This has important implications, since deploying theoretically optimal but complex incentive structures (tax schedules, allocation mechanisms, pricing schemes, etc) may lead to unreliable predictions and suboptimal outcomes — an observation at the root of the ongoing focus on simple mechanisms, both in theoretical (Li2017AER; BorgersLi2019Ecta; PyciaTroyan2023Ecta) and in empirical work (Rees-JonesTaubinsky2020REStud). If problem complexity is taken to be a leading explanation of suboptimal decision-making, understanding which problems are more complex becomes crucial for simplifying them.111 Or to obfuscate them further: a growing literature discusses how complexity can be leveraged and manipulated to the advantage of principals — cf. EllisonWolitzky2012RAND, Spiegler2016ARE, and LiDworczak2022WP.
@@
Response time is often used as an indicator of problem complexity in various settings, including lottery choice (Wilcox1993EJ), loans (KalayciSerra-Garcia2016EE), and portfolio selection (CarvalhoSilverman2017WP).222 Another recent examples are FrydmanNunnari2023WP, who uses decision time together other behavioural patterns in support of specific coordination settings being more complex, and HorensteinGrabiszewski2022JBEE, who use response time to rank extensive-form games in terms of complexity. More examples are surveyed in SpiliopoulosOrtmann2018EE and GillProwse2023EJ. Going beyond economics, a large literature in psychology and neuroscience has also used response time as a proxy for problem complexity (RatcliffMcKoon2008NeCo; ShadlenShohamy2016Neuron; BogaczWagenmakersForstmannNieuwenhuis2010TINS). The underlying assumption is that in more complex problems, making similarly good choices requires greater cognitive effort, and response time can serve as a proxy for cognitive effort (CaplinCsabaLeahyNov2020REStud). If decision-makers are able to make more accurate choices while expending less cognitive effort, as indicated by faster responses, then the decision problem is considered simpler. Therefore, a decision problem is deemed simpler if individuals make fewer mistakes (more accurate choices) and do so with less cognitive effort (faster choices).
@@
The reliance on response time as a proxy for problem complexity is based on the observation that, in many cases, faster choices tend to be more accurate. For example, research in cognitive sciences has indicated that perceptual problems that are harder to discern require greater cognitive effort, leading to slower and less accurate decisions (e.g. ForstmannRatcliffWagenmakers2016ARPsy). Additionally, studies have found that decision-makers take longer when they are closest to being indifferent, indicating the difficulty in distinguishing the best alternative (MostellerNogee1951JPE; KrajbichLuCamererRangel2012FrontiersPsy; Clithero2018JEBO; Alos-FerrerGaragnani2022JRU). This intuition aligns with sequential sampling models (FudenbergStrackStrzalecki2018AER; Goncalves2023WP; Alos-FerrerFehrNetzer2022JPE), where decisions take longer when individuals are closer to being indifferent, as the value of information is highest in such cases. In short, the justification behind the use of response times as a proxy for problem complexity is straightforward: a problem is considered easier if choices are both faster and more accurate, with faster choices often implying greater accuracy.
@@
Whilst response time may be a compelling indicator of problem complexity, it may not always be a reliable measure when used on its own. The existing literature documents a more ambiguous relationship between speed and accuracy, where in some cases faster decisions are better whereas in others slower decisions are more accurate (e.g. AchtzigerAlos-Ferrer2014MnSc; MoritzSiemsenKremer2014POMS; CaplinMartin2016EcIn). Introspection also cautions against positing a monotone relationship between speed, accuracy, and complexity: if problem complexity demands greater cognitive effort and longer decision time, then in excessively complex problems thinking for a long time may not be worthwhile.
@@
In this paper, we reconcile this ambiguous intuition on the relation between speed-accuracy by revisiting a standard Wald problem of optimal stopping (Wald1947Ecta; DvoretzkyKieferWolfowitz1953AMStats). As is conventional in this literature, we take the drift-to-volatility or signal-to-noise ratio as a measure of complexity of the decision problem, capturing the rate of information per unit of time. Our main result is but a simple observation: although choice quality is monotone in problem complexity, expected stopping time is inverse $U$-shaped. This suggests a more nuanced relation between speed and accuracy: When choice problems are simple, people choose fast and well; if they become slightly more complicated, they choose slower and less well; if they become much more complicated, choice quality necessarily deteriorates, but one should instead expect response times to be faster. Extending our model to dynamic effort control, we uncover that this non-monotonicity also suggests restraint in using response times to infer ability of the decision-makers, as high ability individuals always choose better, but do so faster in simple problems and slower in complex ones. Moreover, taking complexity affecting the cost of information acquired and assuming response time is proportional to it, we show that this non-monotonicity is a generic feature of models with costly information acquisition. Finally, we leverage our results to address the crucial question at hand of how to infer problem complexity. We find that distoring incentives is more effective in steering choices in more complex problems, allowing one to infer problem complexity from simple manipulations of incentives.
@@
Our baseline setup is the most canonical sequential sampling model (Wald1947Ecta; DvoretzkyKieferWolfowitz1953AMStats), where a decision-maker faces a binary choice problem with a constant flow cost per unit of time and can obtain information about a binary state of the world by observing a Brownian motion with drift. We take the signal-to-noise ratio, i.e. the drift intensity relative to the instantaneous volatility, as a measure of problem complexity, as is common in existing literature (e.g. Callander2011AER). The nature of the exercise is then to examine how speed (expected stopping time) and accuracy (probability of choosing the optimal alternative) depend on problem complexity.
@@
We then show that, in the canonical setup, accuracy is monotone in problem complexity, whereas expected stopping time is non-monotone. There are two opposing forces that arise when the problem becomes simpler. As signal-to-noise ratio increases, on the one hand, the quality of information is higher, implying that the marginal benefit of learning for longer is greater. On the other hand, the decision-maker can also save on the flow cost of information by stopping earlier. In a standard drift-diffusion version of the model, the decision-maker stops whenever the accumulated evidence exceeds a pre-specified threshold, leading to a monotone relationship between problem complexity, speed, and accuracy.333 This could be interpreted as if the decision-maker is unable to recognise whether the problem is simpler or more complex, and so uses the same threshold.
@@
We leverage closed-form characterisations of the optimal stopping boundaries and expected optimal stopping time to show that the former expand when the problem becomes simpler — entailing greater accuracy — but the latter is non-monotone and single-peaked. This then provides a rationale for the evidence documented by AchtzigerAlos-Ferrer2014MnSc that, in simple situations, quick choices are better, but in more complicated ones, faster decisions are associated with worse performance. Moreover, our results are robust to several variations of the baseline model, including assuming that the drift term is proportional to payoff scale — as in standard drift-diffusion models (FehrRangel2011JEP) — or to having discounting instead of a constant flow cost.
@@
Several papers have examined the relation between response time and ability. For instance, it is often found that slower responses are indicative of higher ability or sophistication (e.g. Rubinstein2016QJE; AgranovCaplinTergiman2015JESA; RecaldeRidelVesterlund2018JPubE; Esteban-CasanellesGoncalves2020WP). However, other studies suggest that, similarly to accuracy, this relationship is ambiguous with greater ability being associated with faster response times in simple problems (DodonovaDodonov2013Intelligence; GoldhammerNaumannStelterTothRolkeKlieme2014JEducPsy; KyllonenZu2016JIntelligence).
@@
We show that this ambiguity is a natural consequence of the non-monotone relation between speed and accuracy. For this we extend our model to allow for the decision-maker to control instantaneous effort, which entails a cost per unit of time. In our model, higher effort scales up the signal-to-noise ratio,444 Our setup is akin to MoscariniSmith2001Ecta, but with no discounting, to keep the model as close as possible to our baseline. and the decision-maker will optimally choose a constant level of effort. This can then be interpreted as if the decision-maker faces an idiosyncratic problem with complexity level that depends on their cost of effort. We contrast the speed and accuracy of high-ability and low-ability decision-makers, where higher ability captures a lower cost of effort and show that high ability individuals always choose better, but do so faster than lower ability individuals in simple problems, and slower in more complex ones. Furthermore, this also cautions against pooling data on multiple individuals, since then response time will not in general be single-peaked.
@@
We then discuss how our results extend to more general models of costly information acquisition. In recent literature it had been pointed out that, in the context of a binary decision problem with one-sided Poisson process, the expected decision time could be non-monotone (BobtcheffLevy2017AEJMicro; HalacKartikLiu2016REStud).555 We note that the non-monotonicity there does seem to depend on modelling details — see BobtcheffLevy2017AEJMicro. Going to a more general setup, we consider general costly information acquisition models (e.g. CaplinDean2015AER), where the decision-maker can acquire a Blackwell experiment to learn about an unknown state paying a cost that is increasing with respect to the Blackwell order. We assume that this cost is scaled up by a function that is increasing in problem complexity.666 In the case of the Wald problem, this is just the square of noise-to-signal ratio (MorrisStrack2019WP). Following evidence from CaplinCsabaLeahyNov2020REStud, we assume that expected stopping time is proportional to the total cost of information. We then show that, in this general setup, expected stopping time is then non-monotone in problem complexity, and provide conditions for it to be single-peaked.
@@
Finally, we return to the fundamental question of how an analyst can infer which problem is more complex. Since response time is not always an appropriate measure of problem complexity, we consider other alternatives. Naturally, when the analyst knows which choice is optimal, they can infer problem complexity from accuracy directly, where the simpler problem is the one in which agents are more likely to choose the best alternative. In many cases, however, which alternative is best depends on unobserved state of the world or idiosyncratic preferences. For instance, the decision problem itself may be meant to have people select different options depending on their preferences or some private type, in which case it is impossible for an analyst to assess if choices are optimal or not.
@@
We propose a simple alternative: to infer problem complexity, the analyst need only to manipulate incentives and observe how choices change. Specifically, we consider the case in which one picks one of the two alternatives in both problems and simply pays decision-makers one extra cent if they choose such alternative. As in any reasonable model, decision-makers would now choose this alternative more often and do so faster, and choose the other less often and slower.777 An observation that holds in for general dynamic costly information acquisition models — see Goncalves2023WP. The analyst can then infer problem complexity from how much more often the incentivised alternatives are now chosen. Specifically, the extra cent has a greater effect on choice probabilities in more complex problems than in simpler ones. In other words, when the problem is simpler, the decision-maker is surer about which alternative is better and so small changes in relative payoffs have a smaller effect on choice probabilities.
@@ 2.  Setup
A decision-maker is choosing from a set $A=\{a,b\}$, with generic element $\alpha\in A$. Payoffs to each alternative depend on an unknown state, $\theta\in\Theta:=\{a,b\}$, and are given by $u:A\times\Theta\to\mathbb{R}$. We assume there are not weakly dominant alternatives so that the choice problem is non-trivial and, without loss of generality, $\theta$ specifies which alternative is best, i.e. $u(\theta,\theta)>u(\theta^{\prime},\theta)$, $\theta\neq\theta^{\prime}$.
We denote the prior probability that $\theta=a$ by $p_{0}\in[0,1]$ and we extend the utility function to distributions over $\Theta$, writing $u(\alpha,p):=\mathbb{E}_{\theta\sim p}[u(\alpha,\theta)]$ and $v(p):=\max_{\alpha}u(\alpha,p)$. For simplicity, we assume each option is equally likely — this simplifies the presentation and the proofs, but does not affect the substance of the results.
@@
Prior to choosing, the decision-maker can acquire information about $\theta$, paying a flow cost per unit of time $c>0$. Formally, the decision-maker observes a continuous-time process given by a Brownian motion with drift, $X_{t}:=\mu s(\theta)t+\sigma B_{t}$, where $B_{t}$ is a standard Brownian motion, $s(\theta)=1$ if $\theta=a$ and $=-1$ if $\theta=b$,888 Note that, given $s(a)\neq s(b)$, insofar as $s$ is known, having the drift being ±$\pm\mu$ is without loss of generality since the process is observationally equivalent. $\mu>0$ corresponds to the drift rate, and $\sigma>0$ scales the instantaneous volatility.
As the signal-to-noise ratio $\mu/\sigma$ captures the rate of information flow per unit of time, we will take this as a measure of problem complexity — with higher values of $\mu/\sigma$ denoting easier problems. This is a common interpretation in economics, cognitive sciences, and psychology (e.g. Callander2011AER; FehrRangel2011JEP; ForstmannRatcliffWagenmakers2016ARPsy), and captures the intuition that in simple problems, with very high $\mu/\sigma$, the optimal choice is apparent or quickly understood.
The model is typically interpreted to be a stylised representation of the process of deliberation or reasoning in decision-making, with the state of the process $X_{t}$ meant to represent accumulated evidence in favour of or against any of the alternatives. While in psychology and neuroscience it became a benchmark in modelling the neural basis of decision-making,999 For instance, it has been the main modelling paradigm for the recalling past experiences and sampling from memory (Ratcliff1978PsyRev; ShadlenShohamy2016Neuron), the sensory perception of physical characteristics (GoldShadlen2001TRENDSCogSci; RatcliffSmithBrownMcKoon2016TrendsCognSci), and brain activation in decision-making (BogaczWagenmakersForstmannNieuwenhuis2010TINS; ShushruthZylberbergShadlen2022CBio). Recent variations extend the model to interpret eye gase, e.g. in purchasing decisions (KrajbichLuCamererRangel2012FrontiersPsy). in economics it has been used to model a wide range of situations and behavioural patterns, including randomness in consumer choice (Clithero2018JEBO; Alos-FerrerFehrNetzer2022JPE), the relation between misperception and small-stakes risk aversion (FrydmanJin2022QJE; KhawLiWoodford2021REStud), consumer’s search for product information (BrancoSunVillas-Boas2012MnSc), and brand recognition (Alos-Ferrer2018JBDM).
@@
The decision-maker then faces a trade-off between taking longer to have a better understanding of which alternative is better, and the effort involved in doing so. Formally, the decision-maker chooses a stopping time $\tau$ adapted to the natural filtration induced by the stochastic process in order to maximise their expected payoff, $\mathbb{E}_{p_{0}}\left[v(p_{\tau})-c\tau\right]$, where $p_{t}$ denotes the posterior belief about $\theta$ given $X_{t}$. That is,
$V(p_{0})=\sup_{\tau}\mathbb{E}_{p_{0}}\left[v(p_{\tau})-c\tau\right].$

### v2

@@ Abstract
This paper re-examines the validity of using response time to infer problem complexity. It revisits a canonical Wald model of optimal stopping, taking signal-to-noise ratio as a measure of problem complexity. While choice quality is monotone in problem complexity, expected stopping time is inverse $U$-shaped. Indeed decisions are fast in both very simple and very complex problems: in simple problems it is quick to understand which alternative is best, while in complex problems it would be too costly — an insight which extends to general costly information acquisition models. This non-monotonicity also underlies an ambiguous relationship between response time and ability, whereby higher ability entails slower decisions in very complex problems, but faster decisions in simple problems. Finally, this paper proposes a new method to correctly infer problem complexity based on the finding that choices react more to changes in incentives in more complex problems.
Keywords: Sequential Sampling; Complexity; Information Acquisition; Drift-Diffusion; Statistical Decision Theory.
JEL Classifications: D83, D84, D90.
@@ 1.  Introduction
It is almost definitional that more complex decision problems entail more mistakes. Individuals make more mistakes and choose dominated alternatives in more complex environments, both in experimental settings (e.g. Rabin and Weizsäcker, 2009; Martínez-Marquina et al., 2019; Banovetz and Oprea, 2023; Enke and Graeber, 2023; Puri, 2023) and in real-world decisions with significant financial implications.111 Examples include choosing health insurance plans (Sinaiko and Hirth, 2011; Bhargava et al., 2017), pension plans (Choi et al., 2011), and mortgages (Keys et al., 2016; Agarwal et al., 2016). This has important implications, since deploying theoretically optimal but complex incentive structures (tax schedules, allocation mechanisms, pricing schemes, etc) may lead to unreliable predictions and suboptimal outcomes — an observation at the root of the ongoing focus on simple mechanisms, both in theoretical and in empirical work.222 For recent theoretical work concerned with simplicity in mechanism design see Li (2017), Börgers and Li (2019), and Pycia and Troyan (2023); an example of empirical research in the same domain is Rees-Jones and Taubinsky (2020). If problem complexity is taken to be a leading explanation of suboptimal decision-making, understanding which problems are more complex becomes crucial for simplifying them.333 Or to obfuscate them further: a growing literature discusses how complexity can be leveraged and manipulated to the advantage of principals — cf. Ellison and Wolitzky (2012), Jehiel (2011), Spiegler (2016), and Li and Dworczak (2022).
@@
Response time is often used as an indicator of problem complexity in various settings, including lottery choice (Wilcox, 1993), loans (Kalaycı and Serra-Garcia, 2016), and portfolio selection (Carvalho and Silverman, 2017).444 Many of these studies show a correlation between response time and problem complexity, rather than assuming it. Other recent examples are Frydman and Nunnari (2023), who use decision time together other behavioural patterns in support of specific coordination settings being more complex, and Horenstein and Grabiszewski (2022), who use it to rank extensive-form games in terms of complexity. More examples are surveyed in Spiliopoulos and Ortmann (2018) and Gill and Prowse (2023). Going beyond economics, a large literature in psychology and neuroscience has also used response time as a proxy for problem complexity (Wright and Ayton, 1988; Ratcliff and McKoon, 2008; Shadlen and Shohamy, 2016; Bogacz et al., 2010). The underlying assumption is that in more complex problems, making similarly good choices requires greater cognitive effort, and response time can serve as a proxy for cognitive effort (Caplin et al., 2020). If decision-makers are able to make more accurate choices while expending less cognitive effort, as indicated by faster responses, then the decision problem is considered simpler. Therefore, a decision problem is deemed simpler if individuals make fewer mistakes (more accurate choices) and do so with less cognitive effort (faster choices).
@@
The reliance on response time as a proxy for problem complexity is based on the observation that, in many cases, faster choices tend to be more accurate. For example, research in cognitive sciences has indicated that perceptual problems that are harder to discern require greater cognitive effort, leading to slower and less accurate decisions (e.g. Forstmann et al., 2016). Additionally, studies have found that decision-makers take longer when they are closest to being indifferent, indicating the difficulty in distinguishing the best alternative (Mosteller and Nogee, 1951; Krajbich et al., 2012; Clithero, 2018; Alós-Ferrer and Garagnani, 2022). This intuition aligns with sequential sampling models (Fudenberg et al., 2018; Gonçalves, 2023; Alós-Ferrer et al., 2021), where decisions take longer when individuals are closer to being indifferent, as the value of information is highest in such cases.555 See also Drugowitsch et al. (2012). In short, the justification behind the use of response times as a proxy for problem complexity is straightforward: a problem is considered easier if choices are both faster and more accurate, with faster choices often implying greater accuracy.
@@
Whilst response time may be a compelling indicator of problem complexity, it may not always be a reliable measure when used on its own. Contributing to an long-standing discussion (Hogarth, 1975; Ford, 1975; Onken et al., 1985; Wright and Ayton, 1988), a more recent literature documents a more ambiguous relationship between speed and accuracy, where in some cases faster decisions are better whereas in others slower decisions are more accurate (e.g. Achtziger and Alós-Ferrer, 2013; Moritz et al., 2014; Caplin and Martin, 2016). Introspection also cautions against positing a monotone relationship between speed, accuracy, and complexity: if problem complexity demands greater cognitive effort and longer decision time, then in excessively complex problems thinking for a long time may not be worthwhile.
@@
In this paper, we reconcile this ambiguous intuition on the relation between speed-accuracy by revisiting a standard Wald problem of optimal stopping (Wald, 1947; Dvoretzky et al., 1953). As is conventional in this literature, we take the drift-to-volatility or signal-to-noise ratio as a measure of complexity of the decision problem, capturing the rate of information per unit of time. Our main result is but a simple observation: although choice quality is monotone in problem complexity, expected stopping time is inverse $U$-shaped. This suggests a more nuanced relation between speed and accuracy: When choice problems are simple, people choose fast and well; if they become slightly more complicated, they choose slower and less well; if they become much more complicated, choice quality necessarily deteriorates, but one should instead expect response times to be faster. Extending our model to dynamic effort control, we uncover that this non-monotonicity also suggests restraint in using response times to infer ability of the decision-makers, as high ability individuals always choose better, but do so faster in simple problems and slower in complex ones. Moreover, taking complexity affecting the cost of information acquired and assuming response time is proportional to it, we show that this non-monotonicity is a generic feature of models with costly information acquisition. Finally, we leverage our results to address the crucial question at hand of how to infer problem complexity. We find that distoring incentives is more effective in steering choices in more complex problems, allowing one to infer problem complexity from simple manipulations of incentives.
@@
Our baseline setup is the most canonical sequential sampling model (Wald, 1947; Dvoretzky et al., 1953), where a decision-maker faces a binary choice problem with a constant flow cost per unit of time and can obtain information about a binary event or parameter by observing a Brownian motion with drift. We take the signal-to-noise ratio, i.e. the drift intensity relative to the instantaneous volatility, as a measure of problem complexity, as is common in existing literature (e.g. Callander, 2011; Fehr and Rangel, 2011). The nature of the exercise is then to examine how speed (expected stopping time) and accuracy (probability of choosing the optimal alternative) depend on problem complexity.
@@
We then show that, in the canonical setup, accuracy is monotone in problem complexity, whereas expected stopping time is non-monotone. While in some cases it may be more important to choose the best alternative than in others, our results show that, no matter relative preference intensity, the probability with which a decision-maker chooses the ex-post optimal alternative monotonically increases as the decision problem becomes simpler. Regarding the stopping time, there are two opposing forces that arise when the problem becomes simpler. As signal-to-noise ratio increases, on the one hand, the quality of information is higher, implying that the marginal benefit of learning for longer is greater. On the other hand, the decision-maker can also save on the cost of information by stopping earlier. In a standard drift-diffusion version of the model, the decision-maker stops whenever the accumulated evidence exceeds an exogenous threshold, leading to a monotone relationship between problem complexity, speed, and accuracy.666 This could be interpreted as if the decision-maker is unable to recognise whether the problem is simpler or more complex, and so uses the same threshold.
@@
We leverage closed-form characterisations of the optimal stopping boundaries and expected optimal stopping time to show that the former expand when the problem becomes simpler — entailing greater accuracy — but the latter is non-monotone and single-peaked. This then provides a rationale for the evidence documented by Achtziger and Alós-Ferrer (2013) that, in simple situations, quick choices are better, but in more complicated ones, faster decisions are associated with worse performance. Moreover, our results are robust to several variations of the baseline model, including assuming that the drift term is proportional to payoff scale — as in standard drift-diffusion models (Fehr and Rangel, 2011) — or to having discounting instead of a constant flow cost.
@@
Several papers have examined the relation between response time and ability. For instance, it is often found that slower responses are indicative of higher ability or sophistication (e.g. Rubinstein, 2016; Agranov et al., 2015; Recalde et al., 2018; Esteban-Casanelles and Gonçalves, 2022). However, other studies suggest that, similarly to accuracy, this relationship is ambiguous, with greater ability being associated with faster response times in simple problems (Dodonova and Dodonov, 2013; Goldhammer et al., 2014; Kyllonen and Zu, 2016). Anecdotal evidence again cautions against generic claims for monotone association: for instance, returning exams earlier typically indicates that a student will score significantly higher or significantly lower than their peers.
@@
We show that this ambiguity is a natural consequence of the non-monotone relation between speed and accuracy. For this we extend our model to allow for the decision-maker to control instantaneous effort, which entails a cost per unit of time. In our model, higher effort scales up the signal-to-noise ratio,777 Our setup is akin to Moscarini and Smith (2001), but with no discounting, to keep the model as close as possible to our baseline. and the decision-maker will optimally choose a constant level of effort. This can then be interpreted as if the decision-maker faces an idiosyncratic problem with complexity level that depends on their cost of effort. We contrast the speed and accuracy of high-ability and low-ability decision-makers, where higher ability captures a lower cost of effort and show that high ability individuals always choose better, but do so faster than lower ability individuals in simple problems, and slower in more complex ones. Furthermore, this also cautions against pooling data on multiple individuals, since then response time will not in general be single-peaked.
@@
We extend our results to more general models of costly information acquisition. Going to a more general setup, we consider general costly information acquisition models (e.g. Caplin and Dean, 2015), where the decision-maker can acquire a Blackwell experiment to learn about an unknown state paying a cost that is increasing with respect to the Blackwell order. We assume that (1) problem complexity corresponds to a parameter that scales up the cost of information,888 As is standard in costly information acquisition or rational inattention models with entropy costs (Matějka and McKay, 2015, e.g.,). and (2) that expected stopping time is proportional to the total cost of information. The first assumption is motivated by the fact that, in the case of the Wald problem, the square of noise-to-signal ratio corresponds to such parameter (Morris and Strack, 2019), whereas the second assumption not only exactly corresponds to our baseline setup, but is also consistent with evidence from Caplin et al. (2020). We then show that, in this general setup, expected stopping time is then non-monotone in problem complexity, and provide conditions for it to be single-peaked.999 While we only consider models in which the cost information is continuous and convex, recent literature has also indicated that in binary decision problem with one-sided Poisson process, the expected decision time could be non-monotone in the arrival rate (Bobtcheff and Levy, 2017; Halac et al., 2016). This non-monotonicity, however, does seem to depend on modelling details — see Bobtcheff and Levy (2017, Proposition 1).
@@
Finally, we return to the fundamental question of how an analyst can infer which problem is more complex. Since response time is not always an appropriate measure of problem complexity, we consider other alternatives. Naturally, when the analyst knows which choice is optimal, they can infer problem complexity from accuracy directly, where the simpler problem is the one in which agents are more likely to choose the best alternative. In many cases, however, which alternative is best depends on unobserved parameter or idiosyncratic preferences. For instance, the decision problem itself may be meant to have people select different options depending on their preferences or some private type, in which case it is impossible for an analyst to assess if choices are optimal or not.
@@
We propose a simple alternative: to infer problem complexity, the analyst need only to manipulate incentives and observe how choices change. Specifically, we consider the case in which one picks one of the two alternatives in both problems and simply pays decision-makers one extra cent if they choose such alternative. As in any reasonable model, decision-makers would now choose this alternative more often and do so faster, and choose the other less often and slower.101010 An observation that holds in for general dynamic costly information acquisition models — see Gonçalves (2023, Theorem 3). The analyst can then infer problem complexity from how much more often the incentivised alternatives are now chosen. Specifically, the extra cent has a greater effect on choice probabilities in more complex problems than in simpler ones. In other words, when the problem is simpler, the decision-maker is surer about which alternative is better and so small changes in relative payoffs have a smaller effect on choice probabilities.
@@ 2.  Setup
A decision-maker is choosing from a set $A=\{a,b\}$, with generic element $\alpha\in A$. Let $(\Omega,\mathcal{F},\mathbb{P})$ denote a probability space, where $\omega\in\Omega$ denotes the state of the world, $\mathcal{F}$ a Borel $\sigma$-algebra, and $\mathbb{P}$ a probability measure. Payoffs to each alternative depend on a parameter or event $\theta\in\Theta:=\{a,b\}$, and are given by $u:A\times\Theta\to\mathbb{R}$. We assume there are not weakly dominant alternatives so that the choice problem is non-trivial and, without loss of generality, $\theta$ specifies which alternative is best, i.e. $u(\theta,\theta)>u(\theta^{\prime},\theta)$, $\theta\neq\theta^{\prime}$.
We denote the prior probability that $\theta=a$ by $p_{0}:=\mathbb{P}(\omega:\theta(\omega)=a)$ and we extend the utility function to distributions over $\Theta$, writing $u(\alpha,p_{0}):=\mathbb{E}[u(\alpha,\theta)]$ and $v(p_{0}):=\max_{\alpha}u(\alpha,p_{0})$ where the dependence on $p_{0}$ is implicit.111111 Instead of understanding varying $p_{0}$ as changing the measure $\mathbb{P}$, it can be equivalently understood as varying the random variable $\theta:\Omega\to\Theta$. For simplicity, we assume each option is equally likely — this simplifies the presentation and the proofs, but does not affect the substance of the results.
@@
Prior to choosing, the decision-maker can acquire information about $\theta$, paying a flow cost per unit of time $c>0$. The decision-maker observes a continuous-time process given by a Brownian motion with drift, $X_{t}:=\mu\,s(\theta)\,t+\sigma\,B_{t}$, where $B_{t}$ is a standard Brownian motion defined on $(\Omega,\mathcal{F},P)$, $s(\theta)=1$ if $\theta=a$ and $=-1$ if $\theta=b$,121212 Note that, given $s(a)\neq s(b)$, insofar as $s$ is known, having the drift being ±$\pm\mu$ is without loss of generality since the process is observationally equivalent. $\mu>0$ corresponds to the drift rate, and $\sigma>0$ scales the instantaneous volatility.
The model is typically interpreted to be a stylised representation of the process of deliberation or reasoning in decision-making, with the state of the process $X_{t}$ meant to represent accumulated evidence in favour of or against any of the alternatives. While in psychology and neuroscience it became a benchmark in modelling the neural basis of decision-making,131313 For instance, it has been the main modelling paradigm for the recalling past experiences and sampling from memory (Ratcliff, 1978; Shadlen and Shohamy, 2016), the sensory perception of physical characteristics (Gold and Shadlen, 2001; Ratcliff et al., 2016), and brain activation in decision-making (Bogacz et al., 2010; Shushruth et al., 2022). Recent variations extend the model to interpret eye gase, e.g. in purchasing decisions (Krajbich et al., 2012). in economics it has been used to model a wide range of situations and behavioural patterns, including randomness in consumer choice (Clithero, 2018; Alós-Ferrer et al., 2021), the relation between misperception and small-stakes risk aversion (Frydman and Jin, 2022; Khaw et al., 2021), consumer’s search for product information (Branco et al., 2012), choosing actions in games (Schotter and Trevino, 2021), information gathering and deliberation in groups (Reshidi et al., 2024), and brand recognition (Alós-Ferrer, 2018).
Problem Complexity
As the signal-to-noise ratio $\mu/\sigma$ captures the rate of information flow per unit of time, we will take this as a measure of problem complexity — with higher values of $\mu/\sigma$ denoting easier problems. This is a common interpretation in economics, cognitive sciences, and psychology (e.g. Callander, 2011; Fehr and Rangel, 2011; Forstmann et al., 2016). In essence, it captures the intuition that simple problems (high $\mu/\sigma$) are those in which a decision-maker quickly realizes which alternative or course of action is best, whereas in complex problems (low $\mu/\sigma$) it would take longer to be as confident. A decision problem may be more complex owing to the intrinsic properties of the items involved — e.g. deciding between savings products is arguably easier and faster than understanding which structured financial product is best for oneself; solving easy exam questions is faster than harder ones. Alternatively, how complex a problem is may have to do with how information is presented or how easy it to obtain information, making it easier or harder to compare across alternatives.
@@
The decision-maker then faces a trade-off between taking longer to have a better understanding of which alternative is better, and the effort involved in doing so. Formally, the decision-maker chooses a stopping time $\tau$ adapted to the natural filtration induced by the stochastic process ${(X_{t})}_{t}$141414 That is, the decision-maker chooses a random variable taking values in $\mathbb{R}_{+}$ such that $\{\omega\in\Omega:\tau(\omega)\leq t\}\in\mathcal{F}_{t}^{X}$, where $\mathcal{F}_{t}^{X}$ is the smallest $\sigma$-field under which $X_{t}$ is measurable. in order to maximise their expected payoff, $\mathbb{E}\left[v(p_{\tau})-c\tau\right]=\mathbb{E}[\max_{\alpha}\mathbb{E}[u(\alpha,\theta)|\mathcal{F}_{\tau}]-c\tau]$, where $p_{t}:=\mathbb{E}[1_{\theta=a}|X_{t}]$ denotes the posterior belief about $\theta$ given $X_{t}$. That is,
$V(p_{0})=\sup_{\tau}\mathbb{E}\left[v(p_{\tau})-c\tau\right],$
where again the dependence on $p_{0}$ is implicit.